# Investigation of the Trailing-Edge Jet-Buffet Phenomenon Using Circulation Control at Transonic Speeds

### v1

@@ 4 Results & Discussion
In circulation control, subsonic studies have shown that the influence of the momentum coefficient on the wing’s lift can be separated from the effects of the angle of attack under pre-stall conditions [34]. This relationship can be expressed as:
$c_{l}=\frac{\partial c_{l}}{\partial\alpha}\Delta\alpha+\frac{\delta c_{l}}{\delta C_{\mu}}\Delta C_{\mu}$ (3)
where $\delta c_{l}/\delta C_{\mu}\approx\Delta c_{l}/C_{\mu}$ represents the augmentation ratio. While the behavior of the augmentation ratio at transonic Mach numbers remains unclear, some studies [35] have provided insights into its correlation with the momentum coefficient. Specifically, for subsonic conditions with low momentum inputs, the augmentation ratio was observed to remain constant when circulation control is maintained. This finding aligns with the trends shown in Fig. 6 (for momentum inputs of $0.003\leq C_{\mu}\leq 0.008$), which illustrates the variation of both the augmentation ratio and lift as functions of the momentum coefficient at an angle of attack of $3^{\circ}$. These results include both experimental data and numerical simulations.
Figure 6: Effect of $C_{\mu}$ on the augmentation ratio and mid-section lift at Mach 0.8 and an angle of attack of $3^{\circ}$.
Where for $\Delta c_{l}\ C_{\mu}$: (   CFD ;   Experiment). And for $c_{l}$: (   CFD   Experiment).
When the momentum input increases beyond a certain threshold, the jet loses its authority, detaches from the surface, and the augmentation ratio rapidly declines to zero with increasing jet strength. For the lift coefficient, a noticeable plateau is observed before this rapid drop occurs. In this region (at momentum inputs of $C_{\mu}\geq 0.008$), attempts to achieve a steady-state solution revealed unsteady behavior in both the aerodynamic coefficients and the residuals. Similar observations were made in other studies [21, 19], although no detailed explanation was provided for this phenomenon.
4.1 Unsteady Flow-Field Description
To evaluate the unsteadiness, 2nd order Unsteady Reynolds-Averaged Navier-Stokes (URANS) simulations were performed at a momentum coefficient of 0.01, corresponding to the first value of the lift plateau. The results were validated through a time step sensitivity study, which considered three different time increments: $\Delta t=(1\times 10^{-2},5\times 10^{-3},1\times 10^{-3})\;\mathrm{(ms)}$, equating to $(128,256,1500)$ iterations per cycle, respectively. Convergence was assessed based on a three-order-of-magnitude reduction in residuals per time step, as well as the stabilization of the lift coefficient’s amplitude and frequency, determined via a one-sided Fast Fourier Transform (FFT). Among the tested time increments, $\Delta t=5\times 10^{-3}\;\mathrm{(ms)}$ was deemed sufficient for convergence, yielding a lift coefficient amplitude of 0.032 and a reduced frequency $\kappa=f_{\mathrm{jet}}r_{e}/\nu_{j}$ of 0.068. Here, $f_{\mathrm{jet}}$ denotes the jet shedding frequency, while $r_{e}$ was approximated as the major radius of the ellipse for simplicity.
Figure 7: Mach fields of the mid-span depicting the vortex shedding evolution from the trailing-edge. With a momentum coefficient of 0.01 at angle of attack of $3^{\circ}$, and free-stream Mach of 0.8.
@@ 4.1 Unsteady Flow-Field Description
An analysis of the flow-field behavior near the trailing-edge indicates that the observed unsteadiness results from periodic detachment of the jet. This phenomenon is illustrated in Fig. 7, where Mach fields are depicted for a portion of a single shedding cycle at various normalized time steps, defined as $\tau=(t-t_{c,0})/T$. Here, $t$ represents the physical flow time, $t_{c,0}$ denotes the initial time of the detachment cycle, and $T$ corresponds to the period. Each image in the figure thus represents the relative temporal position of the Mach fields within the cycle.
@@
Starting from Fig. 7a ($\tau=0$), which corresponds to the maximum lift coefficient ($c_{l}=0.616$), the jet gradually begins to detach from the surface. This progression is observed in Figs. 7(b-d), culminating in full detachment at the minimum lift coefficient ($c_{l}=0.581$ at $\tau=0.23$). Reattachment initiates at $\tau=0.31$, during which the lift coefficient gradually increases (Figs. 7(e-h)), ultimately returning to its peak value, marking the completion of the cycle.
@@
The detachment phase of the cycle is marked by the interaction between the jet and the trailing-edge shockwave, which weakens as the cycle progresses, as well as the interaction between the jet and the underlying boundary-layer. At $\tau=0$, the unsteady field lacks the second pressure peak or separation bubble previously discussed (see Sec. 3) for momentum coefficients at the onset of periodic separation ($C_{\mu}=0.008$). Typically, the separation bubble facilitates mixing between the free-stream momentum and the boundary-layer, thereby delaying detachment [36]. Consequently, its absence can lead to premature separation. In the spatial region where the bubble is present in the steady flow field, the unsteady field instead exhibits a structure formed by the interaction of the shockwave with the boundary-layer. The high-velocity jet pushes the separation bubble closer to the tip of trailing-edge, preventing its stable formation. However, sustaining a periodic cycle of separation and reattachment requires a feedback mechanism. Thus, the unsteady solution alone does not clarify why periodicity is maintained rather than leading to complete jet detachment from the trailing-edge.
@@
RANS computations do not directly resolve the turbulent eddies responsible for mixing, making it difficult to study the "jet-buffet" phenomenon from the perspective of the separation bubble itself. However, previous studies on periodic mechanisms at transonic speeds [27, 28, 29] (see Sec. 1) indicate that the harmonic motion is primarily governed by pressure propagation. Moreover, since circulation control (CC) is predominantly a pressure-driven phenomenon [37], further insights into the jet-buffet can be obtained by analyzing the problem from a pressure-based perspective. As a result, the coupling between a primarily pressure-driven effect and periodic motion suggests that this phenomenon can be effectively analyzed using a dynamics-based approach, such as data-driven methods like Dynamic Mode Decomposition (DMD). DMD extracts the spatial-temporal modes of the flow field while relating them to the system’s dynamic quantities, providing deeper insights into the origins of the phenomenon.
@@ 4.2 Dynamic Mode Decomposition
The modal analysis of the jet-buffet phenomenon is performed using the Dynamic Mode Decomposition (DMD) algorithm. DMD is a data-driven approach for analyzing flow-fields, combining elements of Proper Orthogonal Decomposition (POD) and the Fourier Transform [26]. Like similar techniques, it utilizes Singular Value Decomposition (SVD), a general dimensionality reduction method designed to identify low-dimensional patterns within data [25]. In the context of the observed unsteadiness, DMD serves as a dimensionality reduction tool to capture and investigate the dominant coherent structures that appear to form the jet-buffet. To achieve this, the mean-reduced pressure field is structured into the matrix $X\in\mathbb{R}^{n\times m}$, where the number of rows, $n$, represents the spatial locations of pressure values, and the number of columns, $m$, corresponds to the discrete time steps.
@@
The formal mathematical derivation [25] states that if $X\in\mathbb{C}^{n\times m}$ is a matrix where $n\gg m$, then using Singular Value Decomposition (SVD), $X$ can be decomposed into the left singular matrix, also known as the POD modes, $\Psi\in\mathbb{C}^{n\times n}$, which optimally describes $X$ in a least-squares sense [38], and the right singular matrix $V\in\mathbb{C}^{m\times m}$, both of which are unitary, satisfying $\Psi\Psi^{*}=\Psi^{*}\Psi=I$ and $VV^{*}=V^{*}V=I$, where the superscript $()^{*}$ denotes the conjugate transpose. These matrices satisfy the eigenvalue equation $XV=\Psi\Sigma$, where $\Sigma\in\mathbb{C}^{n\times m}$ is a diagonal matrix containing the data set’s singular values arranged in descending order. To obtain a low-rank approximation, a practical approach involves using the "economy" SVD, which removes the zero-valued elements of $\Sigma$. This leads to the reduced form $X=\hat{\Psi}\hat{\Sigma}\hat{V}^{*}$, where $\hat{\Psi}\in\mathbb{R}^{n\times r}$, $\hat{\Sigma}\in\mathbb{R}^{r\times r}$, $\hat{V}\in\mathbb{R}^{r\times m}$, and $r\leq\min(m,n)$.
@@
The DMD algorithm seeks to determine the matrix $A$ that propagates the snapshots of $X$ forward in time, such that $X^{\prime}\approx AX$, where $X^{\prime}$ represents $X$ advanced by one time step. In a least-squares sense, $A$ is obtained by solving the minimization problem $||X^{\prime}-AX||_{F}$, where $||\cdot||_{F}$ denotes the Frobenius norm. Consequently, the non-linear dynamics of the system are approximated linearly through the matrix $A$. The dynamic modes, or DMD modes, $\Phi$, are the eigenvectors of $A$, associated with the eigenvalues $\Lambda\in\mathbb{C}$, where the diagonal elements $\Lambda_{jj}=\lambda_{j}$ represent the continuous-time eigenvalues. These eigenvalues are mapped as $\lambda=\ln(\mu)/\Delta t$, where $\mu$ denotes the discrete-time eigenvalues of $A$. Using this relation, the system’s linearized modal frequencies and decay rates can be determined by $f=\mathrm{img}(\mathrm{ln}(\mu))/2\pi\Delta t$ and $\xi=\mathrm{real}(\mathrm{ln}(\mu))/\Delta t$ respectively. For a second-order linear system, the natural frequency $f_{n}$ and the damping ratio $\zeta$ are related to $\lambda$ through the equation:
@@
From this, the natural frequency and damping ratio can be computed as $f_{n}=|\lambda|/2\pi$ and $\zeta=-\xi/f_{n}$.
In practice, the modes and frequencies are determined by projecting the matrix $A$ onto the Proper Orthogonal Decomposition (POD) modes [26]. This is expressed as $\tilde{A}=\Psi^{*}A\Psi$, which utilizes the Singular Value Decomposition (SVD), leading to $\tilde{A}=\Psi^{*}X^{\prime}V\Sigma^{-1}$. The matrix $\tilde{A}$ shares the same eigenvalues as $A$ and provides a partial approximation of its eigenvectors. Using the eigendecomposition of $\tilde{A}$, the eigenvectors $W$ are employed to reconstruct the dynamic modes as $\Phi=X^{\prime}V\Sigma^{-1}W$. As a result, the continuous-time solution is expressed as:
$x(t)\approx\sum_{j=1}^{r}\phi_{j}e^{\lambda_{j}t}b_{j}$ (5)
where $b=\Phi^{-1}x(t_{0})$ represents the modal amplitude.
@@ 4.2.1 Preliminary Modal Analysis
The DMD process begins by identifying the dominant modes necessary to accurately capture the jet-buffet phenomenon. By mapping the eigenvalues of $A$ onto the complex plane, as shown in Fig. 8(a), periodicity can be assessed, allowing for the identification of modes associated with the jet-buffet. The data reveals that the eigenvalue distribution on the unit circle contains a cluster of values near $\mathrm{img}(\mu)=0$, where the decay approaches 1 (marked in red in Fig. 8(a)). These eigenvalues correspond to low natural frequencies. Specifically, the eigenvalue $\mu_{\mathrm{jet}}=0.9995+0.0244i$ corresponds to a natural frequency of $f_{\mathrm{jet}}=777.86\;\mathrm{Hz}$ or $\kappa=0.062$, which is consistent with the reduced frequency obtained from the prior FFT analysis. This eigenvalue aligns with the third and fourth modes (ordered by natural frequency) and is under-damped, with a damping ratio of $\zeta_{\mathrm{jet}}=0.00621$, indicating a periodically stable system.
(a) complex mapping of $\mu$.
(b) normalized modal frequencies.
(c) modal amplitude.
(d) modal influence.
Figure 8: Results of the preliminary analysis conducted to ascertain the jet-buffet mode.
Based on this observation, the relationship of each mode to the jet-buffet can be identified by examining the linear dependence of the natural frequencies, normalized by the obtained buffet frequency $f_{\mathrm{jet}}$ (Fig. 8b). Similar studies [28] found that when this dependency breaks, the corresponding modes become disassociated from the dynamic phenomenon. In the present case, the linearity holds for all modes, as no other dynamic effects are present in the accumulated flow-field data. To assess the contribution of these modes to the jet-buffet, a low-rank approximation is made based on their influence on the flow-field, such as through the modal influence [39], which is defined as:
$I_{k}=\sum_{j=1}^{k}b_{j}||\lambda_{j}||$
The modal influence accounts for both the modal amplitude $b$ (depicted independently in Fig. 8c) and the decay rate $\xi$. As shown in Fig. 8(d), capturing 99.9% of the modal influence requires tens to hundreds of modes, in part due to weakly energetic effects that have a significant impact on the dynamics. These effects are given greater weight through the modal influence because of their contribution to the amplitude.
4.2.2 Modal Decomposition

### v3

@@ 4 Results & Discussion
4.1 Unsteady Flow-Field Analysis
The previous section served primarily to establish the physical validity of the results and to reduce numerical uncertainties. The accurate prediction of both global aerodynamic coefficients and pressure distributions at key locations, together with the clear correspondence of the observed phenomenon to experimentally established jet bi-stability, provides strong confidence that the subsequent exploration of the effect rests on firm physical grounds.
@@ 4.1 Unsteady Flow-Field Analysis
Analysis of the bi-stability begins with the investigation of its complete cycle in Fig. 9. The figure presents Mach fields over a portion of a single detachment cycle at various normalized time steps, defined as $\tau=(t-t_{c,0}/t_{c}$. Here, $t$ represents the physical flow time, $t_{c}$ denotes the time of a single cycle and the initial time of the detachment cycle is $t_{c,0}$. Each image in the figure thus represents the relative temporal position of the Mach fields within the cycle.
Figure 9: Mach fields of the mid-span depicting the jet bi-stability. With a momentum coefficient of 0.01 at angle of attack of $3^{\circ}$, and free-stream Mach of 0.8.
@@
Starting from Fig. 9a ($\tau=0$), which corresponds to the maximum lift coefficient ($c_{l}=0.616$), the jet gradually begins to detach from the surface. This progression is observed in Figs. 9(b-d), culminating in full detachment at the minimum lift coefficient ($c_{l}=0.581$ at $\tau=0.23$). Reattachment initiates at $\tau=0.31$, during which the lift coefficient gradually increases (Figs. 9(e-h)), ultimately returning to its peak value, marking the completion of the cycle ($\tau\geq 0.5$).
@@
A key feature of this phenomenon is the interaction between the jet and the trailing-edge shockwave, which weakens as the cycle progresses, as well as the interaction between the jet and the underlying boundary-layer. At $\tau=0$, and as previously observed in the ensemble data, the unsteady field lacks the pressure peak or separation bubble previously discussed (see Sec. 3) for momentum coefficients at the onset of periodic separation ($C_{\mu}\geq 0.008$). Typically, the separation bubble facilitates mixing between the free-stream momentum and the boundary-layer, thereby delaying detachment [36]. Consequently, its absence can lead to premature separation. The high-velocity pushes the separation bubble closer to the tip of trailing-edge, preventing its stable formation.
(a) wing
(b) trailing-edge
Figure 10: Pressure coefficient distribution at mid-span for different parts of the cycle, at a free-stream Mach number of 0.8, angle of attack of $3^{\circ}$ and a momentum coefficient of 0.01.
$\tau=0$,   $\tau=0.15$ ,   $\tau=0.31$,   $\tau=0.5$
To further delineate on the observation, the pressure data explicitly analyzed in Fig. 10 which depicts the pressure distribution at various stages of the cycle. In Fig. 10(b), the absence of the bubble at $\tau=0$ manifests as a sharp, linear drop in the pressure coefficient within the range $0.99\leq x/c\leq 1$. As the cycle progresses, this drop becomes more gradual, as seen at $\tau=0.15$. This change in slope arises from the jet’s inability to sustain a stable separation bubble given the current trailing-edge length, thus supporting the initial detachment caused by the downstream displacement of the bubble. A potential approach to mitigate this would be to extend the trailing-edge length, thereby promoting the formation of a stable bubble. However, this comes at the expense of a flatter pressure distribution curve, which reduces lift enhancement. At $\tau=0.31$, when the reattachment phase begins, the pressure distribution near $x/c=0.985$ gradually becomes positive, indicating an adverse pressure gradient acting on the flow, a characteristic signature of flow separation. This unfavorable gradient diminishes as the cycle reaches $\tau=0.5$, where suction is re-established and a relatively flat pressure distribution occurs, corresponding to reduced lift and the attaining of bi-stability. Despite clear modifications to the pressure field near the trailing-edge caused by the periodic detachment, the upstream surface pressure distribution, depicted in Fig. 10(a) remains relatively constant throughout the cycle, reinforcing the decoupling of the bi-stability from the upper-surface shockwave.
@@
The unsteady solution, although underlines the initial detachment as the incapability of the flow to maintain a stable bubble, does not clarify why periodicity is maintained rather than leading to complete jet detachment from the trailing-edge. Furthermore RANS computations do not directly resolve the turbulent eddies responsible for mixing, making it difficult to study the bi-stability phenomenon from the perspective of the separation bubble itself. However, previous studies on periodic mechanisms at transonic speeds [29, 30, 31] (see Sec. 1) indicate that the harmonic motion is primarily governed by pressure propagation. Moreover, since circulation control (CC) is predominantly a pressure-driven phenomenon [37], further insights into the bi-stability can be obtained by analyzing the problem from a pressure-based perspective. As a result of the pressure dominance, the coupling between a primarily pressure-driven effect and periodic motion suggests that this phenomenon can be effectively analyzed using a dynamics-based approach, such as data-driven methods like Dynamic Mode Decomposition (DMD). DMD extracts the spatial-temporal modes of the flow field while relating them to the system’s dynamic quantities. In this context DMD can provide deeper insights into the phenomenon.
@@ 4.2 Dynamic Mode Decomposition
The modal analysis of the bi-stability is performed using the Dynamic Mode Decomposition (DMD) algorithm. DMD is a data-driven approach for analyzing flow-fields, combining elements of Proper Orthogonal Decomposition (POD) and the Fourier Transform [28]. Like similar techniques, it utilizes Singular Value Decomposition (SVD), a general dimensionality reduction method designed to identify low-dimensional patterns within data [27]. In the context of the observed unsteadiness, DMD serves as a dimensionality reduction tool to capture and investigate the dominant coherent structures that appear to form the bi-stability as well as whether it can be represented in a reduced form. To achieve this, the mean-reduced pressure field is structured into the matrix $X\in\mathbb{R}^{n\times m}$, where the number of rows, $n$, represents the spatial locations of pressure values, and the number of columns, $m$, corresponds to the discrete time steps.
@@
The formal mathematical derivation [27] states that if $X\in\mathbb{C}^{n\times m}$ is a matrix where $n\gg m$, then using Singular Value Decomposition (SVD), $X$ can be decomposed into the left singular matrix, also known as the POD modes, $\Psi\in\mathbb{C}^{n\times n}$, which optimally describes $X$ in a least-squares sense [38], and the right singular matrix $V\in\mathbb{C}^{m\times m}$, both of which are unitary, satisfying $\Psi\Psi^{*}=\Psi^{*}\Psi=I_{\mathrm{id}}$ and $VV^{*}=V^{*}V=I_{\mathrm{id}}$, where the superscript $()^{*}$ denotes the conjugate transpose and $I_{\mathrm{id}}$ is the identity matrix. These matrices satisfy the eigenvalue equation $XV=\Psi\Sigma$, where $\Sigma\in\mathbb{C}^{n\times m}$ is a diagonal matrix containing the data set’s singular values arranged in descending order. To obtain a low-rank approximation, a practical approach involves using the "economy" SVD, which removes the zero-valued elements of $\Sigma$. This leads to the reduced form $X=\hat{\Psi}\hat{\Sigma}\hat{V}^{*}$, where $\hat{\Psi}\in\mathbb{R}^{n\times r}$, $\hat{\Sigma}\in\mathbb{R}^{r\times r}$, $\hat{V}\in\mathbb{R}^{r\times m}$, and $r\leq\min(m,n)$.
@@
The DMD algorithm seeks to determine the matrix $A$ that propagates the snapshots of $X$ forward in time, such that $X^{\prime}\approx AX$, where $X^{\prime}$ represents $X$ advanced by one time step. In a least-squares sense, $A$ is obtained by solving the minimization problem $X^{\prime}-AX||_{F}$, where $||\cdot||_{F}$ denotes the Frobenius norm. Consequently, the non-linear dynamics of the system are approximated linearly through the matrix $A$. The spatial distributions $\Phi$, are the eigenvectors of $A$, associated with the eigenvalues $\Lambda\in\mathbb{C}$, where the diagonal elements $\Lambda_{jj}=\lambda_{j}$ represent the continuous-time eigenvalues. These eigenvalues are mapped as $\lambda=\ln(\mu)/\Delta t$, where $\mu$ denotes the discrete-time eigenvalues of $A$. Using this relation, the system’s linearized modal frequencies and decay rates can be determined by
$f=\mathrm{img}(\mathrm{ln}(\mu))/2\pi\Delta t$ (3)
$\xi=\mathrm{real}(\mathrm{ln}(\mu))/\Delta t$ (4)
respectively. For a second-order linear system, the natural frequency $f_{n}$ and the damping ratio $\zeta$ are related to $\lambda$ through the equation:
@@
From this, the natural frequency and damping ratio can be computed as $f_{n}=|\lambda|/2\pi$ and $\zeta=-\xi/f_{n}$. In practice, the modes and frequencies are determined by projecting the matrix $A$ onto the Proper Orthogonal Decomposition (POD) modes [28]. This is expressed as $\tilde{A}=\Psi^{*}A\Psi$, which utilizes the Singular Value Decomposition (SVD), leading to $\tilde{A}=\Psi^{*}X^{\prime}V\Sigma^{-1}$. The matrix $\tilde{A}$ shares the same eigenvalues as $A$ and provides a partial approximation of its eigenvectors. Using the eigendecomposition of $\tilde{A}$, the eigenvectors $W$ are employed to reconstruct the dynamic modes as $\Phi=X^{\prime}V\Sigma^{-1}W$. As a result, the continuous-time solution for the pressure field is expressed as:
$p(t)\approx\sum_{j=1}^{r}\phi_{j}e^{\lambda_{j}t}b_{j}$ (6)
where $b=\Phi^{-1}p(t(0))$ represents the modal amplitude.
@@ 4.2.1 Preliminary Modal Analysis
Figure 11: Analysis of the pressure field mode scatter behavior for the modal mapping to the complex plane and their transformation to dynamic properties.
[…]
The DMD process begins by identifying the dominant modes that capture the bi-stability phenomenon. Fig. 11 shows the eigenvalues of matrix $A$ plotted on the complex plane for the total 267 snapshots used, provides insight into the dynamic stability of each mode. The vicinity of the eigenvalues to the unit circle is an indicator of the mode’s dynamic stability. Modes that are naturally stable tend closer to the unit circle [28], while those further from the circle are more unstable and contribute less to the sustained dynamics. To further interpret these modes, they are transformed in terms of their damping ratio $\zeta$ and reduced frequency $\kappa$ (see Sec. 4.1). Red marked points, where $\zeta<0$, represent modes that amplify over time, destabilizing the system due to amplification of the exponential term in Eq.6 and are referred throughout as the "driving modes", while blue marked points with $0\leq\zeta\leq 1$ indicate "underdamped modes", that decay or oscillate stably. Because the bi-stability is characterized by a low-frequency periodic motion, the relevant mode is expected to lie at the minimum reduced frequency, i.e., $\kappa_{\mathrm{jet}}=\kappa_{\mathrm{min}}$, with a damping ratio tending toward zero. In Fig. 11(a), the eigenvalue $\mu_{\mathrm{jet}}=0.9995+0.0244i$ corresponds to a frequency of $f_{\mathrm{jet}}=777.86\,\mathrm{Hz}$ or $\kappa=0.096$, meeting these criteria. This mode is marked with a black circle and aligns well with the dominant frequency identified in the prior FFT analysis.
To further assess their relevance, the contribution of these modes to the bi-stability, is examined based on their influence on the flow-field. Although a variety of parameters which analyze accumulating contributions are present in the literature, this study utilizes the modal influence [39]. It accounts for both the modal amplitude $b$ and the decay rate $\xi$ adding weight to weakly energetic effects through their contribution to the exponential term. The modal influence, $I$ is defined as
[…]
Figure 12: Contribution of the accumulating modal influence relative to the total influence with increasing mode number.
[…]
Fig. 12 illustrates the modal normalized influence, $I/I_{\mathrm{max}}$ as a function of mode number, while Fig. 13 shows the amplitude of both. As observed in Fig. 12, the first mode ($r=1$), corresponding to the time-averaged pressure field as well as the second mode ($r=2$), exhibit almost no influence on the bi-stability, unlike subsequent modes. Notably, 24.5% of the total modal influence is concentrated in less than 2% of the underdamped modes, namely the two modal pairs at $r=3,4$ and $r=5,6$. Although a full reconstruction of the flow field would require more modes, these pairs contribution is significant to the dynamics.
Figure 13: Contribution of the modal amplitude with increasing mode number.
[…]
In Fig. 13, the amplitude of each mode is shown to decay with increasing mode number. The dominant underdamped modes up to the second modal pair ($r=4,5$) display an order-of-magnitude difference in amplitude compared with later modes. The driving modes, while present, are of lower magnitude and exhibit more erratic behavior. A clear contrast in trends is seen between the underdamped and driving modes, with the former following a more consistent decay. It is thus observed that the impact of the first modal pairs is substantial. Consequently, constructing a reduced model of the bi-stability requires examining these modes directly in relation to the rest and ascertaining their influence on the pressure field.
4.2.2 Modal Spatial Distribution